# Energy exchange between electrons and ions driven by ITG-TEM turbulence

### v1

@@ III.2 Turbulent energy exchange in pure TEM turbulence
[…]
Figure 3 shows all terms in Eq. (9) in a steady state of turbulence. For clarity, each term for electrons and ions is normalized by $q^{\mathrm{turb}}_{e}/(T_{e}L_{Te})$ and $q^{\mathrm{turb}}_{e}/(T_{i}L_{Te})$, respectively. The numerical error in the entropy balance, defined as the difference between the left- and right-hand sides of Eq. (9) is within 12 % of $q^{\mathrm{turb}}_{e}/(T_{e}L_{Te})$ for electrons and 3 % for ions, while $Q^{\mathrm{turb}}_{i}=-Q^{\mathrm{turb}}_{e}$ is 6 %. This error arises from the numerical diffusion as well as from losses due to outflow at the $z$-coordinate boundaries. Although these issues can be mitigated by increasing the resolution in the $z$-coordinate and extending the $z$-domain, the resolution and domain employed in this study are sufficient for evaluating fluxes and energy exchanges.
@@
This situation is analogous to that in ITG turbulence in terms of transferring energy from particle species with large entropy production due to turbulent transports to particle species with small entropy production. In TEM turbulence, it is found that electrons generate larger entropy through turbulent transport than ions, and the direction of energy exchange is from electrons to ions. Thus, it is confirmed that the conjecture regarding the direction of turbulent energy transfer, as discussed in Sec. II, is applicable to pure TEM turbulence as well.
@@
The wavenumber spectrum of each component of turbulent energy transfer in Eq. (10) is shown in Fig. 4. First, the collisional components $Q^{\mathrm{turb}}_{aC}(a=e,i)$ are found to have negligible influence on the turbulent energy exchange. Similarly, the nonlinear interaction components between different wavenumbers, $Q^{\mathrm{turb}}_{a\psi}(a=e,i)$, have small impact across the wavenumber spectrum. As in ITG turbulence, both parallel and perpendicular heating terms, $Q^{\mathrm{turb}}_{a\parallel}$ and $Q^{\mathrm{turb}}_{aB}(a=e,i)$, predominantly determine the magnitude and direction of energy transfer in TEM turbulence. However, the magnitudes of these contributions for electrons and ions differ from those in ITG turbulence. The turbulent energy transfer in ITG turbulence is mainly composed of perpendicular cooling for ions $Q^{\mathrm{turb}}_{iB}<0$ and parallel heating for electrons $Q^{\mathrm{turb}}_{e\parallel}>0$. In contrast, the energy transfer in TEM turbulence is dominated by perpendicular cooling for electrons $Q^{\mathrm{turb}}_{eB}<0$ and parallel heating for ions $Q^{\mathrm{turb}}_{i\parallel}>0$. These results suggest that the signs of parallel heating $Q^{\rm turb}_{a\parallel\bm{k}}>0\ (a=e,i)$ and perpendicular cooling $Q^{\rm turb}_{aB\bm{k}}<0\ (a=e,i)$ except for zonal mode ($k_{y}=0$) remain unchanged when the dominant instability changes. This indicates that the direction of turbulent energy exchange is determined by the relative magnitudes of these components.
Next, we investigate the applicability of quasilinear modeling for turbulent energy exchange in pure TEM turbulence. The quasilinear model estimates the nonlinear simulation results based on two components: (i) the ratio of fluxes or energy exchange to the squared amplitude of the electrostatic potential, obtained from linear simulations, and (ii) the saturation rule of the squared amplitude of the electrostatic potential in a steady state turbulence. In this study, we focus on the former and do not deal with the latter. The ratio is calculated by
[…]
Figures 5 (a) and (b) show the wavenumber spectra of terms in the entropy balance equation in Eq. (1) in pure TEM turbulence. Entropy produced by particle and heat fluxes for electrons is partially transferred to ions. Additionally, entropy is redistributed among wavenumber modes via nonlinear interactions $N_{e\bm{k}}$ and $N_{i\bm{k}}$, which transfer entropy from unstable modes to zonal and high-wavenumber modes. Subsequently, entropy is dissipated through collisions in a wide wavenumber range, which maintains the overall entropy balance. This mechanism is similar to the situation in ITG turbulence, except that the roles of electrons and ions are reversed.
Figures 5 (c) and (d) show the wavenumber spectra of nonlinear and quasilinear weights in pure TEM turbulence. The nonlinear weights calculated at $k_{x}\rho_{ti}=0$, $WX^{0}_{a,N}(X_{a}=\Gamma_{a},q_{a},Y_{a}$ for $a=e,i)$ are found to agree well with those obtained by summing over $k_{x}$ modes, $WX_{a,N}$, respectively, which indicates that the result from linear simulations for $k_{x}=0$ mode is appropriate for evaluating quasilinear weights. It is found that quasilinear weights for energy exchange sufficiently agree with their nonlinear weights within an error margin of $15\%$ or less in the colored wavenumber region where $N_{e\bm{k}}<0$ and $N_{i\bm{k}}<0$ , while those for particle and heat fluxes are within a factor of two except for ion heat flux. Furthermore, the discrepancy between quasilinear and nonlinear weights becomes more significant in higher wavenumber modes, beyond the colored wavenumber region. Since the colored wavenumber region accounts for $91\%$ of the total energy exchange, the predictability of the quasilinear model for energy exchange in pure TEM turbulence is validated.
@@ III.3 Turbulent energy exchange in mixed ITG-TEM turbulence
Figure 7: Turbulent energy transfer in mixed ITG-TEM turbulence as a function of $\Delta EP$ (a), $\Delta FP$ (b), and $\Delta EF$ (c), as defined in Eqs. (34)-(36), respectively. The physical dimensional quantities are calculated by using $\delta\equiv\rho_{ti}/R_{0}=1.4\times 10^{-3},$ $n_{e}=1.0\times 10^{20}\ \mathrm{m^{-3}},$ $T_{i}=2.0\ \mathrm{keV}$.
@@
Figure 10: For electrons and ions $(a=e,i)$, the parallel heating $Q^{\rm turb}_{a\parallel}$ (left) and perpendicular cooling $Q^{\rm turb}_{aB}$ (right) are plotted as functions of $(\mathcal{E}^{\mathrm{turb}}_{e}+\mathcal{E}^{\mathrm{turb}}_{i})/2R_{0}$, and $\mathcal{E}^{\mathrm{turb}}_{a}/R_{0}$, respectively.
@@
are also used on the horizontal axis for comparison in Figs. 7 (b) and (c), respectively. As illustrated in Fig. 7 (a), the sign of energy exchange is consistent with that of $\Delta EP$. Therefore, this investigation indicates that energy is transferred from a particle species with larger entropy production due to turbulent transports to the other species in mixed ITG-TEM turbulence, and the validity of the conjecture is confirmed. Furthermore, the signs of $\Delta FP$ and $\Delta EF$ are also found to coincide with that of the energy exchange, suggesting a linear correlation between energy exchange and them. Since the slope of energy transfer as a function of $\Delta EF$ is approximately unity, this result supports the validity of Eq. (24).
@@
Figure 7 suggests the correlation between energy flux and energy exchange as indicated in Eq. (24). Then, we examine the validity of the hypothesis presented in Eqs. (22) and (25). Figure 10 shows the parallel heating and perpendicular cooling components of energy exchange for electrons and ions as functions of $(\mathcal{E}^{\mathrm{turb}}_{e}+\mathcal{E}^{\mathrm{turb}}_{i})/2R_{0}$, and $\mathcal{E}^{\mathrm{turb}}_{a}/R_{0}$, respectively. It is found that the perpendicular coolings for electrons and ions are proportional to each energy flux. In particular, they appear to be on the same straight line passing through the origin, and the slope is approximately -1. Then, the validity of Eq. (22) is demonstrated. Although the parallel heating exhibits greater scatter than the perpendicular heating, it maintains a proportional relationship, with a slope of unity, to the average energy flux of each particle species divided by the major radius. From these results, it is clarified that the perpendicular heating terms are primarily determined by the energy flux of their respective particle species.
@@
Finally, the feasibility of the quasilinear model for energy exchange in ITG-TEM turbulence is examined. Figures 11 and 12 show the wavenumber spectra of terms in the entropy balance equation along with nonlinear and quasilinear weights for $(R_{0}/L_{Te},R_{0}/L_{Ti})=(5.0,3.0)$ and $(4.0,4.0)$, respectively. These two cases are regarded as the results of turbulence where both ITG and TEM coexist. In the case of $(R_{0}/L_{Te},R_{0}/L_{Ti})=(5.0,3.0)$, entropy is largely produced by the electron particle and heat fluxes, and entropy transfer from electrons to ions is the main source of entropy production for ions. This situation is similar to pure TEM turbulence. For both electrons and ions $(a=e,i)$, the values of nonlinear weights for energy transfer calculated at $k_{x}\rho_{ti}=0$, $WY^{0}_{a,N}$, are approximately half the values of $WY_{a,N}$ in the wavenumber regions where $N_{a\bm{k}}<0$. The quasilinear weights for energy exchange $WY^{0}_{a,L}$ agree with $WY^{0}_{a,N}$ in the wavenumber regions. Although the discrepancy between the nonlinear and quasilinear weights for energy transfer is larger than that observed in pure ITG or TEM turbulence, the level of agreement remains comparable to that of particle and heat fluxes within a factor of two. A similar trend is observed for all parameter sets except for $(R_{0}/L_{Te},R_{0}/L_{Ti})=(4.0,4.0)$.
@@
where $\mathcal{E}_{a,QL}=q_{a,QL}+5T_{a}\Gamma_{a,QL}/2$. The results calculated by Eq.(38), along with $\Delta EF$ computed from the nonlinear simulation, are shown in Fig.6. The energy flux difference between electrons and ions from the nonlinear simulation $\Delta EF$ shows good agreement with the turbulent energy exchange, including the sign. These results suggest that the model based on Eq. (24) or (36) has the potential to predict turbulent energy exchange effectively. It is found that the signs of energy exchange obtained by Eq. (38) match those from nonlinear simulations for all parameter sets, including the critical case of $(R_{0}/L_{Te},R_{0}/L_{Ti})=(4.0,4.0)$. On the other hand, because the quasilinear weights for particle and heat fluxes tend to be overestimated, the magnitudes of energy exchange obtained from Eq. (38) are roughly twice as large as the nonlinear simulation results. This discrepancy can be mitigated by employing a model that more accurately predicts energy fluxes or by applying an appropriate correction factor.
@@ IV Conclusions and Discussion
In this study, the turbulent energy transfer in pure TEM and mixed ITG-TEM turbulence in tokamak plasmas is investigated. It is confirmed that pure TEM turbulence transfers energy from electrons to ions, while pure ITG turbulence transfers energy from ions to electrons. The wavenumber spectral analysis clarifies that energy is transferred from electrons to the perturbed electrostatic potential via the cooling of electrons in the $\nabla B$-curvature drift motion and the heating of ions streaming along the field line, which are the main mechanisms of turbulent energy exchange in pure TEM turbulence. From the viewpoint of entropy balance, the particle and heat fluxes of electrons generate large entropy associated with the perturbed distribution function, and energy exchange plays a role in transferring part of the entropy to ions. Thus, TEM turbulence is expected to facilitate the transfer of energy from alpha-heated electrons to ions, contributing to more efficient ion heating in fusion plasmas.
The investigation of mixed ITG-TEM turbulence reveals that the direction of energy transfer reverses as the most unstable mode transitions between ITG and TEM. It is also confirmed that the sign of energy exchange aligns with that of the difference in entropy production. Therefore, the conjecture that energy is transferred by turbulence from particle species with large entropy production due to turbulent transports to the other particle species is found to be valid.
@@
In order to examine the predictability of the quasilinear model in ITG-TEM turbulence, quasilinear weights for turbulent fluxes and energy exchange are also investigated. In the case of turbulence where the unstable mode is significantly biased toward ITG or TEM, the quasilinear weight for energy exchange agrees with the nonlinear weight within an error margin of 20%, which is better than that for particle and heat fluxes. On the other hand, when both ITG and TEM instabilities contribute comparably to the turbulence, the accuracy of quasilinear predictions for turbulent fluxes remains almost unchanged, whereas that for energy exchange deteriorates. In addition, it is revealed that when the wavenumber spectrum of energy transfer contains both positive and negative contributions, the quasilinear model may predict the opposite sign of energy exchange compared to the result obtained from nonlinear simulations. In such cases, it is worthwhile to examine quasilinear weights by taking into account not only the most unstable mode but also the subdominant unstable modesStaebler2021 ; Staebler2024 ; Merz .

### v2

@@ III.2 Turbulent energy exchange in pure TEM turbulence
where $\langle\cdots\rangle_{s}$ denotes the surface average without ensemble or time averaging. The time evolution of instantaneous energy transfer $\breve{Q}^{\rm turb}_{a}$ in pure TEM turbulence is shown in Fig. 2. The instantaneous energy transfer associated with ions $\breve{Q}^{\rm turb}_{i}$ exhibits larger spike-like structures compared to that of electrons $\breve{Q}^{\rm turb}_{e}$, and the instantaneous net turbulent heating $\sum_{a=e,i}\breve{Q}^{\rm turb}_{a}$ is not necessarily zero at every moment in time. Over a longer timescale in a steady state, however, the sum approaches zero, indicating that there is no net energy generation and that energy is being exchanged between electrons and ions. The time average in a steady state of turbulence is used instead of the ensemble average to evaluate $\langle\langle\cdots\rangle\rangle$. In the steady state, the time-averaged turbulent heatings $Q^{\rm turb}_{e}$ and $Q^{\rm turb}_{i}$ are 180-180 and $184~[\delta^{2}n_{e}T_{i}v_{ti}/R_{0}]$, respectively, with standard deviations of $84$ and $210~[\delta^{2}n_{e}T_{i}v_{ti}/R_{0}]$. Thus, in the magnetic configuration of CBC, energy is transferred from electrons to ions in TEM turbulence, which is in the direction opposite to that in ITG turbulence. Therefore, TEM turbulence in tokamak palasmas is expected to facilitate energy transfer from alpha-heated electrons to ions, potentially enhancing the efficiency of ion heating.
Figure 3 shows all terms in Eq. (10) in a steady state of turbulence. For clarity, each term for electrons and ions is normalized by $q^{\mathrm{turb}}_{e}/(T_{e}L_{Te})$ and $q^{\mathrm{turb}}_{e}/(T_{i}L_{Te})$, respectively. The numerical error in the entropy balance, defined as the difference between the left- and right-hand sides of Eq. (10) is within 12 % of $q^{\mathrm{turb}}_{e}/(T_{e}L_{Te})$ for electrons and 3 % for ions, while $Q^{\mathrm{turb}}_{i}=-Q^{\mathrm{turb}}_{e}$ is 6 %. This error arises from the numerical diffusion as well as from losses due to outflow at the $z$-coordinate boundaries. When the resolution in the $z$-coordinate is doubled, the error associated with the numerical diffusion is reduced by 1/16. Although these issues can be mitigated by increasing the resolution in the $z$-coordinate and extending the $z$-domain, the resolution and domain employed in this study are sufficient for evaluating fluxes and energy exchangesKato2024 .
@@
This situation is analogous to that in ITG turbulence in terms of transferring energy from particle species with large entropy production due to turbulent transport to particle species with small entropy production. In TEM turbulence, it is found that electrons generate larger entropy through turbulent transport than ions, and the direction of energy exchange is from electrons to ions. Thus, it is confirmed that the conjecture regarding the direction of turbulent energy transfer, as discussed in Sec. II, is applicable to pure TEM turbulence as well.
@@
The wavenumber spectrum of each component of turbulent energy transfer in Eq. (12) is shown in Fig. 4. First, the collisional components $Q^{\mathrm{turb}}_{aC}(a=e,i)$ are found to have negligible influence on the turbulent energy exchange. Similarly, the nonlinear interaction components between different wavenumbers, $Q^{\mathrm{turb}}_{a\psi}(a=e,i)$, have small impact across the wavenumber spectrum. As in ITG turbulence, both parallel and perpendicular heating terms, $Q^{\mathrm{turb}}_{a\parallel}$ and $Q^{\mathrm{turb}}_{aB}(a=e,i)$, predominantly determine the magnitude and direction of energy transfer in TEM turbulence. However, the magnitudes of these contributions for electrons and ions differ from those in ITG turbulence. The turbulent energy transfer in ITG turbulence is mainly composed of perpendicular cooling for ions $Q^{\mathrm{turb}}_{iB}<0$ and parallel heating for electrons $Q^{\mathrm{turb}}_{e\parallel}>0$. In contrast, the energy transfer in TEM turbulence is dominated by perpendicular cooling for electrons $Q^{\mathrm{turb}}_{eB}<0$ and parallel heating for ions $Q^{\mathrm{turb}}_{i\parallel}>0$, as also clearly shown in Fig. 8. These results suggest that the signs of parallel heating $Q^{\rm turb}_{a\parallel\bm{k}}>0\ (a=e,i)$ and perpendicular cooling $Q^{\rm turb}_{aB\bm{k}}<0\ (a=e,i)$ except for zonal mode ($k_{y}=0$) remain unchanged when the dominant instability changes. This indicates that the direction of turbulent energy exchange is determined by the relative magnitudes of these components.
@@ III.3 Turbulent energy exchange in mixed ITG-TEM turbulence
Figure 6: Turbulent energy transfer in mixed ITG-TEM turbulence as a function of $\Delta EP$ (a), $\Delta FP$ (b), and $\Delta EF$ (c), as defined in Eqs. (36)-(38), respectively. The physical dimensional quantities are calculated by using $\delta\equiv\rho_{ti}/R_{0}=1.4\times 10^{-3},$ $n_{e}=1.0\times 10^{20}\ \mathrm{m^{-3}},$ $T_{i}=2.0\ \mathrm{keV}$. The dot line in Fig. 6(c) represents $Q^{\rm turb}_{i}=(\mathcal{E}^{\rm turb}_{e}-\mathcal{E}^{\rm turb}_{i})/(2R_{0})$.
@@
Figure 9: For electrons and ions $(a=e,i)$, the parallel heating $Q^{\rm turb}_{a\parallel}$ (left) and perpendicular cooling $Q^{\rm turb}_{aB}$ (right) are plotted as functions of $(\mathcal{E}^{\mathrm{turb}}_{e}+\mathcal{E}^{\mathrm{turb}}_{i})/2R_{0}$, and $\mathcal{E}^{\mathrm{turb}}_{a}/R_{0}$, respectively. The dot lines represent $Q^{\rm turb}_{a\parallel}=(\mathcal{E}^{\rm turb}_{e}+\mathcal{E}^{\rm turb}_{i})/(2R_{0})$ and $Q^{\rm turb}_{aB}=-\mathcal{E}^{\rm turb}_{a}/R_{0}$, respectively.
@@
are also used on the horizontal axis for comparison in Figs. 6 (b) and (c), respectively. As illustrated in Fig. 6 (a), the sign of energy exchange is consistent with that of $\Delta EP$. Therefore, this investigation indicates that energy is transferred from a particle species with larger entropy production due to turbulent transport to the other species in mixed ITG-TEM turbulence, and the validity of the conjecture is confirmed. Furthermore, the signs of $\Delta FP$ and $\Delta EF$ are also found to coincide with that of the energy exchange, suggesting a linear correlation between energy exchange and them. Since the slope of energy transfer as a function of $\Delta EF$ is approximately unity, this result supports the validity of Eq. (26).
@@
Having verified the correlation between energy flux and energy exchange defined in Eq. (26), we now examine the validity of the hypothesis presented in Eqs. (24) and (27). Figure 9 shows the parallel heating and perpendicular cooling components of energy exchange for electrons and ions as functions of $(\mathcal{E}^{\mathrm{turb}}_{e}+\mathcal{E}^{\mathrm{turb}}_{i})/2R_{0}$, and $\mathcal{E}^{\mathrm{turb}}_{a}/R_{0}$, respectively. It is found that the perpendicular coolings for electrons and ions are proportional to each energy flux. In particular, they appear to be on the same straight line passing through the origin, and the slope is approximately -1. Then, the validity of Eq. (24) is demonstrated. Although the parallel heating exhibits greater scatter than the perpendicular heating, it maintains a proportional relationship, with a slope of unity, to the average energy flux of each particle species divided by the major radius. From these results, it is clarified that the perpendicular heating terms are primarily determined by the energy flux of their respective particle species.
III.4 Quasilinear model for ITG-TEM turbulence
@@
Finally, we investigate the applicability of quasilinear model for turbulent energy exchange in ITG-TEM turbulence. Figures 10 (a) and (b) show the wavenumber spectra of terms in the entropy balance equation in Eq. (3) in pure TEM turbulence. Entropy produced by particle and heat fluxes for electrons is partially transferred to ions. Additionally, entropy is redistributed among wavenumber modes via nonlinear interactions $N_{e\bm{k}}$ and $N_{i\bm{k}}$, which transfer entropy from unstable modes to zonal and high-wavenumber modes. Although the entropy transfer from unstable to stable modes within the same $k_{y}$ region may also occur under this condition Terry , such an analysis has not been conducted in this study. Meanwhile, entropy is dissipated through collisions in a wide wavenumber range, which maintains the overall entropy balance. This mechanism is similar to the situation in ITG turbulence, except that the roles of electrons and ions are reversed.
Figures 10 (c) and (d) show the wavenumber spectra of nonlinear and quasilinear weights in pure TEM turbulence. The nonlinear weights calculated at $k_{x}\rho_{ti}=0$, $WX^{0}_{a,N}(X_{a}=\Gamma_{a},q_{a},Y_{a}$ for $a=e,i)$ are found to agree well with those obtained by summing over $k_{x}$ modes, $WX_{a,N}$, respectively, which indicates that the result from linear simulations for $k_{x}=0$ mode is appropriate for evaluating quasilinear weights. It is found that quasilinear weights for energy exchange sufficiently agree with their nonlinear weights within an error margin of $15\%$ or less in the colored wavenumber region where $N_{e\bm{k}}<0$ and $N_{i\bm{k}}<0$ , while those for particle and heat fluxes are within a factor of two except for ion heat flux. Furthermore, the discrepancy between quasilinear and nonlinear weights becomes more significant in higher wavenumber modes, beyond the colored wavenumber region. Since the colored wavenumber region accounts for $91\%$ of the total energy exchange, the validity of the quasilinear model for energy exchange in pure TEM turbulence is confirmed.
Next, the feasibility of the quasilinear model for energy exchange in ITG-TEM turbulence is examined. Figures 11 and 12 show the wavenumber spectra of terms in the entropy balance equation along with nonlinear and quasilinear weights for $(R_{0}/L_{Te},R_{0}/L_{Ti})=(5.0,3.0)$ and $(4.0,4.0)$, respectively. These two cases are regarded as the results of turbulence where both ITG and TEM coexist. In the case of $(R_{0}/L_{Te},R_{0}/L_{Ti})=(5.0,3.0)$, entropy is largely produced by the electron particle and heat fluxes, and entropy transfer from electrons to ions is the main source of entropy production for ions. This situation is similar to pure TEM turbulence. For both electrons and ions $(a=e,i)$, the values of nonlinear weights for energy transfer calculated at $k_{x}\rho_{ti}=0$, $WY^{0}_{a,N}$, are approximately half the values of $WY_{a,N}$ in the wavenumber regions where $N_{a\bm{k}}<0$. The quasilinear weights for energy exchange $WY^{0}_{a,L}$ agree with $WY^{0}_{a,N}$ in the wavenumber regions. Although the discrepancy between the nonlinear and quasilinear weights for energy transfer is larger than that observed in pure ITG or TEM turbulence, the level of agreement remains comparable to that of particle and heat fluxes within a factor of two. A similar trend is observed for all parameter sets except for $(R_{0}/L_{Te},R_{0}/L_{Ti})=(4.0,4.0)$.
@@
where $\mathcal{E}_{a,QL}=q_{a,QL}+5T_{a}\Gamma_{a,QL}/2$. The results calculated by Eq.(40), along with $\Delta EF$ computed from the nonlinear simulation, are shown in Fig.5. The energy flux difference between electrons and ions from the nonlinear simulation $\Delta EF$ shows good agreement with the turbulent energy exchange, including the sign. These results suggest that the model based on Eq. (26) or (38) has the potential to predict turbulent energy exchange effectively. It is found that the signs of energy exchange obtained by Eq. (40) match those from nonlinear simulations for all parameter sets, including the critical case of $(R_{0}/L_{Te},R_{0}/L_{Ti})=(4.0,4.0)$. On the other hand, because the quasilinear weights for particle and heat fluxes tend to be overestimated, the magnitudes of energy exchange obtained from Eq. (40) are roughly twice as large as the nonlinear simulation results. This discrepancy can be mitigated by employing a model that more accurately predicts energy fluxes or by applying an appropriate correction factor. However, the signs do not necessarily match perfectly even with Eq. (40). In such cases, the magnitude of the turbulent energy exchange is very small and can therefore be neglected. Hence, this sign inconsistency is acceptable.
@@ IV Conclusions and Discussion
In this study, the turbulent energy transfer in pure TEM and mixed ITG-TEM turbulence in tokamak plasmas is investigated. It is confirmed that pure TEM turbulence transfers energy from electrons to ions, while pure ITG turbulence transfers energy from ions to electrons. The wavenumber spectral analysis clarifies that energy is transferred from electrons to the perturbed electrostatic potential via the cooling of electrons in the $\nabla B$-curvature drift motion and from the potential to ions via heating of ions streaming along the field line, which are the main mechanisms of turbulent energy exchange in pure TEM turbulence. From the viewpoint of entropy balance, the particle and heat fluxes of electrons generate large entropy associated with the perturbed distribution function, and energy exchange plays a role in transferring part of the entropy to ions. Thus, TEM turbulence is expected to facilitate the transfer of energy from alpha-heated electrons to ions, contributing to more efficient ion heating in fusion plasmas.
The investigation of mixed ITG-TEM turbulence reveals that the direction of energy transfer reverses as the most unstable mode transitions between ITG and TEM. It is also confirmed that the sign of energy exchange aligns with that of the difference in entropy production. Therefore, this result supports the conjecture that energy is transferred by turbulence from particle species with large entropy production due to turbulent transport to the other species.
@@
In order to examine the validity of the quasilinear model in ITG-TEM turbulence, quasilinear weights for turbulent fluxes and energy exchange are also investigated. In the case of turbulence where the unstable mode is significantly biased toward ITG or TEM, the quasilinear weight for energy exchange agrees with the nonlinear weight within an error margin of 20%, which is better than that for particle and heat fluxes. On the other hand, when both ITG and TEM instabilities contribute comparably to the turbulence, the accuracy of quasilinear predictions for turbulent fluxes remains almost unchanged, whereas that for energy exchange deteriorates. In addition, it is revealed that when the wavenumber spectrum of energy transfer contains both positive and negative contributions, the quasilinear model may predict the opposite sign of energy exchange compared to the result obtained from nonlinear simulations. In such cases, it is worthwhile to examine quasilinear weights by taking into account not only the most unstable mode but also the subdominant unstable modesStaebler2021 ; Staebler2024 ; Merz .